OpenStudy (anonymous):

A stone is thrown vertically upward with a speed of 50m/s from the edge of a cliff that is 275 m high. 1)how much later in seconds does it reach the bottom of the cliff? 2) what is the total distance in meters the stone travels before hitting the ground?

OpenStudy (anonymous):

I have the answers if that helps, I just don't know how to get them

OpenStudy (anonymous):

1) is 14 seconds 2) x=530

OpenStudy (anonymous):

\(h(t)=\frac{a}{2}t^2+vt+h_0\) where \(a\)= -9.8 m/s² or -32 ft/s², depending on the units \(v\) = initial velocity \(h_0\) = initial height So your equation is \[h(t)=-4.9t^2+50t+275\] Set that equal to 0 and solve for t and you'll get 14.2 s.

OpenStudy (anonymous):

Where did you get 4.9?

OpenStudy (anonymous):

-9.8/2 = -4.9 because your length is in meters

OpenStudy (anonymous):

I was told to use either vf=vi+at Or Vf^2=vi+2a × delta x

OpenStudy (anonymous):

neither of those helps you. You're not looking for and you don't have final velocities. In your notation the formula I used would be \[\Delta x=v_i+\frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

Okaythanks

OpenStudy (anonymous):

sorry forgot a t \[\Delta x=v_i t+\frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

How about 2?

OpenStudy (anonymous):

You can use the second formula you put up for part 2. The velocity is 0 at the max height, so \(v_f=0\) \[v_f^2=v_i^2+2ad\] \[0=(50 ~m/s)^2+2(-9.8~m/s^2)d\] d = 127 m so it goes 127 m on the way up. It goes (127 m + 275 m) on the way down. Total = 530.1 m

OpenStudy (anonymous):

Omg,thanks so much! Now I understand a little better

OpenStudy (anonymous):

you're welcome